# Three-dimensional fracton topological orders with boundary Toeplitz braiding

### v4

@@ III.3 Condition for Type-I and Type-II iCS field theories
Combining (51) and (62), we conclude that if the elements of the $K$-matrix satisfy
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3},\end{array}\right.$ (67)
the $K$-matrix exhibits two boundary zero modes, and the corresponding iCS field theory is a Type-I theory. Furthermore, combining the condition (36), (44) and (57), we conclude that if the elements of $K$-matrix satisfy
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\\ l_{2}\neq l_{3},\end{array}\right.\text{ or }\left\{\begin{array}[]{l}|m|<|l_{2}-l_{3}|,\\ l_{4}\neq 0,\\ \det B=l_{1}l_{4}-l_{2}l_{3}=0,\\ (2l_{2}m-l_{4}n)(2l_{3}m-l_{4}n)=0,\end{array}\right.$ (76)
the $K$-matrix possesses only 1 boundary zero zero mode, and the corresponding iCS field theory is a Type-II iCS field theory. The aforementioned $K_{1}^{2N}$ matrix, in its mathematical form, is equivalent to the single-particle Hamiltonian of a typical 1d topological insulator i.e., the SSH model. Since $K_{1}^{2N}$ has two boundary zero modes, it belongs to Type-I iCS field theory. Generally speaking, Type-I $K$-matrix has been widely discussed in the study of TI and TSC. The presence of boundary zero mode, as is well-known, can be indicated by topological invariants. However, the Toeplitz matrices with only one boundary zero mode, i.e., the Type-II Toeplitz $K$-matrices, are rarely studied in the context of TI and TSC. Although we draw comparisons between the $K$-matrices of iCS field theories and the Hamiltonians of TI and TSC, it is important to note that the occurrence of Toeplitz braiding does not require any symmetry protection.
@@ IV Numerical computation on iCS Field Theories with Toeplitz braiding
In Section IV.1 and IV.2, we conduct a numerical computation that illustrates the connection between the appearance of nonlocal braiding statistics in $z$-direction and the existence of boundary zero modes of $K$-matrix. For comparative analysis, we focus on iCS field theories that arise from stacking $\mathbb{Z}_{2}$ twisted topological orders. The block $A$ in the $K$-matrix, given by (2), is $\begin{pmatrix}2&2\\ 2&0\end{pmatrix}$, which describes the double-semion topological order. In Section IV.3, we present the symmetry fractionalization pattern in iCS field theories enriched by $U(1)$ charge conservation symmetry.
@@ IV.1 Numerical computation on Toeplitz braiding via Type-I and Type-II iCS field theories
To demonstrate the correspondence between boundary zero modes and Toeplitz braiding in Type-I and Type-II iCS field theories, we consider three iCS field theories as illustrative examples without loss of generality. The $K$-matrix (Eq. (2)) are denoted as $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$, where $2N$ is the system size as well as the size of the $K$-matrix. Block $A$ and block $B$ in $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$ matrices are denoted as $A_{j}$ and $B_{j}$ ($j=2,3,4$), where
@@
The elements of $K_{2}^{2N}$ satisfy the condition (62) in Sec. III, implying that $K_{2}^{2N}$ has two boundary zero modes. Therefore, the corresponding iCS field theory is a Type-I iCS field theory. The elements of $K_{3}^{2N}$ satisfy the condition (36) , whereas $K_{4}^{2N}$ satisfy the condition (44). Therefore, $K_{3}^{2N}$ possess one lower boundary zero mode while $K_{4}^{2N}$ exhibit one upper boundary zero modes. They belong to Type-II iCS field theories. To demonstrate the nonlocal braiding statistics encoded in these iCS field theories, i.e., the Toeplitz braiding, we numerically compute the braiding statistics encoded in finite-size $K$-matrices of different dimensions, from which how the braiding statistics scale with system size is known. The mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ is presented in Fig. 5, where $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ (Eq. (16a)) is the mutual braiding phase angle between topological excitations carrying unit gauge charge (Eq. (6)). Fig. 5(a1-a3), Fig. 5(b1-b3) and Fig. 5(c1-c3) are the matrix plots of mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ of $K_{2}$-matrix, $K_{3}$-matrix and $K_{4}$-matrix iCS field theories, respectively, where we set the size of the $K$-matrix to be $20$, $30$ and $40$. Numerical results show the robust presence of non-zero elements in the upper-right and lower-left positions of the matrices upon the system size is enlarged, which is vital to Toeplitz braiding. Fig. 5(a4) shows that how the mutual braiding phase angle $\Theta_{\mathbf{l}_{2},\mathbf{l}_{2N}}$ changes with respect to the system size $2N$, while Fig. 5(b4) and Fig. 5(c4) show how the mutual braiding phase angle $\Theta_{\mathbf{l}_{1},\mathbf{l}_{2N}}$ changes with respect to the system size $2N$. From the numerical results, we clearly find that the phase angle oscillates as the system size $2N$ increases, indicative of the existence of Toeplitz braiding.
@@
To more clearly understand the nontriviality of Toeplitz braiding, it is beneficial to make a sharp comparison by demonstrating what we expect for the $K$-matrix that doesn’t support Toeplitz braiding. For this purpose, we consider a new $K$-matrix denoted as $K_{5}^{2N}$, composed of the following $A$ and $B$ blocks
@@ IV.3 Charge fractionalization in iCS field theory
In contrast, for the (gapped) iCS field theories without Toeplitz braiding, even with OBC in $z$, $Q_{I}$ exhibits translational symmetry for $I$ corresponding to the $z$-bulk coordinates ($1\ll I\ll 2N$), such that $Q_{I}\simeq Q_{I+2}$, consistent with the charge fractionalization pattern observed under PBC. In Fig. 7(a5) and Fig. 7(b5), we present the fractionalized charges $Q_{I}$ of topological excitations in $K_{5}$-matrix iCS field theories. Fig. 7(a5) is under OBC while Fig. 7(b5) is under PBC. From this, we understand that in the absence of boundary zero modes of the $K$-matrix, (gapped) iCS field theories with PBC provide a good description to the $z$-bulk physics in OBC.
@@ V Conclusions
There are several future directions. First of all, the categorization of iCS field theories with Toeplitz braiding is primarily for mathematical convenience. How to physically distinguish Type-I and Type-II iCS field theories is left for future research. It will also be interesting to construct the lattice model realization of iCS field theories that supports boundary Toeplitz braiding, which may pave the way for further numerical or experimental study on Toeplitz braiding. From the perspective of fractional quantum Hall states or spin liquid, the stacking process described in Sec. II.1 can be realized by condensing topological excitations followed by adding interlayer interaction of condensation currents or spin-spin interaction [44, 45, 15, 12], which may be also useful here. Detailed construction that explicitly enables Toeplitz braiding will be left to future discussion. We might also consider enlarging blocks $A$ and $B$ in the Toeplitz $K$-matrix, which might affect the number of boundary zero modes and reshape the braiding statistics. Breaking the strict Toeplitz character of the $K$-matrix and considering the “stability” of braiding statistics against stacking faults is also an intriguing topic. Furthermore, following the construction of topologically-ordered non-liquid states in this paper, the construction of higher-dimensional topologically-ordered non-liquid states is directly allowed. For instance, by stacking (3+1)D topological orders [46, 47, 48, 49], we can obtain a (4+1)D topologically-ordered non-liquid state. Similarly, stacking (2+1)D topological orders in two independent spatial dimensions can also yield (4+1)D topologically-ordered non-liquid states. The study of these theories and the duality between them is left for future research. We can also compute entanglement entropy via partitioning the system into two subsystems along $z$-direction and study whether and how the presence of Toeplitz braiding leaves its fingerprint in the scaling behavior of entanglement entropy.
@@ Acknowledgements.
Appendix A Detailed Derivation on Sorting Boundary Zero Modes of Toeplitz $K$-matrices
The Appendix provides further discussions on the technical details of sorting boundary zero modes of the $K$-matrix in Sec. III of the main text. To make reading easier, we list the key equations from the main text below: A trial solution $\mathbf{u}$ of boundary zero mode is a linear combination of boundary bases $\{\mathbf{w}^{j}\}$:
@@ Appendix A Detailed Derivation on Sorting Boundary Zero Modes of Toeplitz $K$-matrices
A boundary base $\mathbf{w}^{j}$ has the form

### v5

@@ III.3 Condition for Type-I and Type-II iCS field theories
Taking the union of (39) and (82), we find that the $K$ matrix Eq. (2) has a boundary zero mode matching the boundary condition Eq. (16a) if one of the following conditions is satisfied:
$\mathbb{M}_{1}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{3}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{M}_{2}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|,\\ \end{array}\right.\quad\text{or}\quad\mathbb{M}_{3}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|.\end{array}\right.$ (110)
Similarly, taking the union of (50) and (90), we find that the $K$ matrix Eq. (2) has a boundary zero mode matching the boundary condition Eq. (16b) if one of the following conditions is satisfied:
$\mathbb{N}_{1}:~{}\left\{\begin{array}[]{l}m\neq 0,\quad l_{4}\neq 0,\\ l_{1}l_{4}=l_{2}l_{3},\quad 2l_{2}m=l_{4}n,\\ |m|<|l_{3}-l_{2}|,\end{array}\right.\quad\text{or}\quad\mathbb{N}_{2}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\quad n(l_{2}+l_{3})=2l_{1}m,\\ |m|<|l_{2}+l_{3}|,\quad|l_{2}|<|l_{3}|,\\ \end{array}\right.\quad\text{or}\quad\mathbb{N}_{3}:\left\{\begin{array}[]{l}m\neq 0,\\ l_{4}=0,\\ |m|<|l_{2}+l_{3}|,\quad|l_{3}|<|l_{2}|.\end{array}\right.$ (120)
The set operation $(\mathbb{M}_{1}\cup\mathbb{M}_{2}\cup\mathbb{M}_{3})\cap(\mathbb{N}_{1}\cup\mathbb{N}_{2}\cup\mathbb{N}_{3})$ renders that the $K$ matrix possesses two boundary zero modes if the following condition is satisfied:
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})=2l_{1}m,\\ l_{2}\neq l_{3}.\end{array}\right.$ (125)
In this case, the corresponding iCS field theory is a Type-I theory. Furthermore, the set operation $(\mathbb{M}_{1}\cup\mathbb{M}_{2}\cup\mathbb{M}_{3}\cup\mathbb{N}_{1}\cup\mathbb{N}_{2}\cup\mathbb{N}_{3})\backslash[(\mathbb{M}_{1}\cup\mathbb{M}_{2}\cup\mathbb{M}_{3})\cap(\mathbb{N}_{1}\cup\mathbb{N}_{2}\cup\mathbb{N}_{3})]$ renders the following conditions for exactly one boundary zero mode:
$\left\{\begin{array}[]{l}|m|<|l_{2}+l_{3}|,\\ l_{4}=0,\\ n(l_{2}+l_{3})\neq 2l_{1}m,\\ l_{2}\neq l_{3},\end{array}\right.\text{ or }\left\{\begin{array}[]{l}|m|<|l_{2}-l_{3}|,\\ l_{4}\neq 0,\\ \det B=l_{1}l_{4}-l_{2}l_{3}=0,\\ (2l_{2}m-l_{4}n)(2l_{3}m-l_{4}n)=0.\end{array}\right.$ (134)
In this case, the corresponding iCS field theory is a Type-II iCS field theory. More specifically, if $\left\{\begin{array}[]{l}l_{4}=0,\\ |l_{2}|<|l_{3}|,\end{array}\right.$ or $\left\{\begin{array}[]{l}l_{4}\neq 0,\\ 2l_{3}m=l_{4}n,\end{array}\right.$ the $K$ matrix has a lower boundary zero mode satisfying the boundary condition Eq. (16a); if $\left\{\begin{array}[]{l}l_{4}=0,\\ |l_{3}|<|l_{2}|,\end{array}\right.$ or $\left\{\begin{array}[]{l}l_{4}\neq 0,\\ 2l_{2}m=l_{4}n,\end{array}\right.$ the $K$ matrix has an upper boundary zero mode satisfying the boundary condition Eq. (16b). The results are summarized in Table 1. The aforementioned $K_{1}^{2N}$ matrix, in its mathematical form, is equivalent to the single-particle Hamiltonian of a typical 1d topological insulator i.e., the SSH model. Since $K_{1}^{2N}$ has two boundary zero modes, it belongs to Type-I iCS field theory. Generally speaking, Type-I $K$ matrix has been widely discussed in the study of TI and TSC. The presence of boundary zero mode, as is well-known, can be indicated by topological invariants. However, the Toeplitz matrices with only one boundary zero mode, i.e., the Type-II Toeplitz $K$ matrices, are rarely studied in the context of TI and TSC. Although we draw comparisons between the $K$ matrices of iCS field theories and the Hamiltonians of TI and TSC, it is important to note that the occurrence of Toeplitz braiding does not require any symmetry protection.
@@ IV Numerical computation on iCS Field Theories with Toeplitz braiding
In Section IV.1 and IV.2, we conduct a numerical computation that illustrates the connection between the appearance of nonlocal braiding statistics in $z$-direction and the existence of boundary zero modes of $K$ matrix. For comparative analysis, we focus on iCS field theories that arise from stacking $\mathbb{Z}_{2}$ twisted topological orders. The block $A$ in the $K$ matrix, given by (2), is $\begin{pmatrix}2&2\\ 2&0\end{pmatrix}$, which describes the double-semion topological order. In Section IV.3, we present the symmetry fractionalization pattern in iCS field theories enriched by $U(1)$ charge conservation symmetry.
@@ IV.1 Numerical computation on Toeplitz braiding via Type-I and Type-II iCS field theories
To demonstrate the correspondence between boundary zero modes and Toeplitz braiding in Type-I and Type-II iCS field theories, we consider three iCS field theories as illustrative examples without loss of generality. The $K$ matrices (Eq. (2)) are denoted as $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$, where $2N$ is the system size as well as the size of the $K$ matrix. Block $A$ and block $B$ in $K_{2}^{2N}$, $K_{3}^{2N}$ and $K_{4}^{2N}$ matrices are denoted as $A_{j}$ and $B_{j}$ ($j=2,3,4$), where
@@
The elements of $K_{2}^{2N}$ satisfy the condition (95) in Sec. III, implying that $K_{2}^{2N}$ has two boundary zero modes. Therefore, the corresponding iCS field theory is a Type-I iCS field theory. The elements of $K_{3}^{2N}$ satisfy the condition (39) , whereas $K_{4}^{2N}$ satisfy the condition (50). Therefore, $K_{3}^{2N}$ possess one lower boundary zero mode while $K_{4}^{2N}$ exhibit one upper boundary zero modes. They belong to Type-II iCS field theories. To demonstrate the nonlocal braiding statistics encoded in these iCS field theories, i.e., the Toeplitz braiding, we numerically compute the braiding statistics encoded in finite-size $K$ matrices of different dimensions, from which how the braiding statistics scale with system size is known. The mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ is presented in Fig. 5, where $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ (Eq. (16a)) is the mutual braiding phase angle between topological excitations carrying unit gauge charge (Eq. (6)). Fig. 5(a1-a3), Fig. 5(b1-b3) and Fig. 5(c1-c3) are the matrix plots of mutual braiding phase angle $\Theta_{\mathbf{l}_{I},\mathbf{l}_{J}}$ of $K_{2}$-matrix, $K_{3}$-matrix and $K_{4}$-matrix iCS field theories, respectively, where we set the size of the $K$ matrix to be $20$, $30$ and $40$. Numerical results show the robust presence of non-zero elements in the upper-right and lower-left positions of the matrices upon the system size is enlarged, which is vital to Toeplitz braiding. Fig. 5(a4) shows that how the mutual braiding phase angle $\Theta_{\mathbf{l}_{2},\mathbf{l}_{2N}}$ changes with respect to the system size $2N$, while Fig. 5(b4) and Fig. 5(c4) show how the mutual braiding phase angle $\Theta_{\mathbf{l}_{1},\mathbf{l}_{2N}}$ changes with respect to the system size $2N$. From the numerical results, we clearly find that the phase angle oscillates as the system size $2N$ increases, indicative of the existence of Toeplitz braiding.
@@
To more clearly understand the nontriviality of Toeplitz braiding, it is beneficial to make a sharp comparison by demonstrating what we expect for the $K$ matrix that doesn’t support Toeplitz braiding. For this purpose, we consider a new $K$ matrix denoted as $K_{5}^{2N}$, composed of the following $A$ and $B$ blocks
@@ IV.3 Charge fractionalization in iCS field theory
In contrast, for the (gapped) iCS field theories without Toeplitz braiding, even with OBC in $z$, $Q_{I}$ exhibits translational symmetry for $I$ corresponding to the $z$-bulk coordinates ($1\ll I\ll 2N$), such that $Q_{I}\simeq Q_{I+2}$, consistent with the charge fractionalization pattern observed under PBC. In Fig. 7(a5) and Fig. 7(b5), we present the fractionalized charges $Q_{I}$ of topological excitations in $K_{5}$-matrix iCS field theories. Fig. 7(a5) is under OBC while Fig. 7(b5) is under PBC. From this, we understand that in the absence of boundary zero modes of the $K$ matrix, (gapped) iCS field theories with PBC provide a good description to the $z$-bulk physics in OBC.
@@ V Conclusions
There are several future directions. First of all, the categorization of iCS field theories with Toeplitz braiding is primarily for mathematical convenience. How to physically distinguish Type-I and Type-II iCS field theories is left for future research. It will also be interesting to construct the lattice model realization of iCS field theories that supports boundary Toeplitz braiding, which may pave the way for further numerical or experimental study on Toeplitz braiding. From the perspective of fractional quantum Hall states or spin liquid, the stacking process described in Sec. II.1 can be realized by condensing topological excitations followed by adding interlayer interaction of condensation currents or spin-spin interaction [45, 46, 16, 12], which may be also useful here. Detailed construction that explicitly enables Toeplitz braiding will be left to future discussion. We might also consider enlarging blocks $A$ and $B$ in the Toeplitz $K$ matrix, which might affect the number of boundary zero modes and reshape the braiding statistics. Breaking the strict Toeplitz character of the $K$ matrix and considering the “stability” of braiding statistics against stacking faults is also an intriguing topic. Furthermore, following the construction of topologically-ordered non-liquid states in this paper, the construction of higher-dimensional topologically-ordered non-liquid states is directly allowed. For instance, by stacking (3+1)D topological orders [47, 48, 49, 50], we can obtain a (4+1)D topologically-ordered non-liquid state. Similarly, stacking (2+1)D topological orders in two independent spatial dimensions can also yield (4+1)D topologically-ordered non-liquid states. The study of these theories and the duality between them is left for future research. We can also compute entanglement entropy via partitioning the system into two subsystems along $z$-direction and study whether and how the presence of Toeplitz braiding leaves its fingerprint in the scaling behavior of entanglement entropy.
@@ Acknowledgements.
Appendix A Detailed Derivation on Sorting Boundary Zero Modes of Toeplitz $K$ matrices
The Appendix provides further discussions on the technical details of sorting boundary zero modes of the $K$ matrix (Eq. (2)) in Sec. III of the main text. To facilitate reading, we outline the sorting procedures below, and list key equations. As is described in Sec. III, the ansatz for a trial solution $\mathbf{u}$ of boundary zero mode satisfying $K\mathbf{u}=0$ in the thermodynamic limit $N\rightarrow\infty$ is a linear combination of boundary bases $\{\mathbf{w}^{j}\}$:
@@ Appendix A Detailed Derivation on Sorting Boundary Zero Modes of Toeplitz $K$ matrices
Here the boundary base $\mathbf{w}^{j}$ has the form
@@
The solutions $\{\beta_{j}\}$ to Eq. (146) together with the bulk recurrence relation Eq. (144) settle down the boundary bases $\{\mathbf{w}^{j}\}$ in Eq. (143) (up to a normalization factor). The boundary bases are utilized to construct trial solutions. It is required that a trial solution $\mathbf{u}$ localizes at only one $z$-boundary, i.e., all the $\{\beta_{j}\}$ in one trial solution $\mathbf{u}$ satisfies either $|\beta_{j}|>1$ or $|\beta_{j}|<1$. If a trial solution $\mathbf{u}$ is indeed a legitimate boundary zero mode consisting of boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|<1$, it satisfies the boundary condition
$A\begin{pmatrix}u_{1}^{a}\\ u_{1}^{b}\end{pmatrix}+B^{T}\begin{pmatrix}u_{2}^{a}\\ u_{2}^{b}\end{pmatrix}=0.$ (151a)
Likewise, if a trial solution $\mathbf{u}$ is a legitimate boundary zero mode consisting of boundary bases $\{\mathbf{w}^{j}\}$ with $|\beta_{j}|>1$, it satisfies the boundary condition
Therefore, we can identify the conditions on the elements of the $K$ matrix that allow trial solutions to satisfy the boundary condition in Eq. (151a) or Eq. (151b), i.e. the conditions that enable boundary zero modes. As is discussed in the main text, we classify our discussion based on the degree of Eq. (146), which affects the # of boundary bases as well as the structure of trial solutions. More concretely, the degree of Eq. (146) is determined by $\alpha_{1}=l_{1}l_{4}-l_{2}l_{3}=\det B$.
If $\det B=0$, Eq. (146) reduces to a cubic equation. In this case, one solution is $\beta_{0}=0$, which is discarded, leaving the two remaining solutions $\beta_{1}$ and $\beta_{2}$. They are determined by the quadratic equation